PRESERVATION OF THE ABSOLUTELY CONTINUOUS SPECTRUM OF SCHRÖDINGER EQUATION UNDER

PERTURBATIONS BY SLOWLY DECREASING POTENTIALS AND A.E. CONVERGENCE OF INTEGRAL OPERATORS

ALEXANDER KISELEV

Abstract. We prove new criteria of stability of the absolutely continuous spectrum of one-dimensional Schrödinger operators under slowly decaying perturbations. As applications, we show that the absolutely continuous spectrum of the free and periodic Schrödinger operators is preserved under perturbations by all potentials V (x) satisfying |V (x)| ≤ C(1 + x)^{− \frac{2}{3} −ϵ}. The main new technique includes an a.e. convergence theorem for a class of integral operators.